OpenStudy (anonymous):

which best describes the discriminant of the graph shown? the graph is an upside down parabola

OpenStudy (owlcoffee):

The discriminant of a quadratic function stems from the general formula: \[ax^2+bx+c=0 \iff x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] The discriminant is what determines if the function will intersect the x-axis one, two or not intersect it at all. It is simbolized with the capital delta, and equals all the coefficients inside the square root, which are \(b^2-4ac\) and it allows us to study the functions behaviour without having to actually use the whole formula: \[\Delta = b^2-4ac\]